# Learning with Calibration: Exploring Test-Time Computing of Spatio-Temporal Forecasting

### v1

@@ E.1 Complete Results Table
w/ ST-TTC $34.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $52.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.54}}$ $29.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.79}}$ $41.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.62}}$ $63.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $27.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.03}}$
$\Delta$downarrow 1.3% $\downarrow$ 1.2% $\downarrow$ 12.5% $\downarrow$ 3.9% $\downarrow$ 4.4% $\downarrow$ 9.1%
@@
Avg. STONE $28.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $41.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.26}}$ $36.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.91}}$ $33.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.69}}$ $49.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.97}}$ $29.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.00}}$
@@
w/ ST-TTC $26.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.22}}$ $39.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $24.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.00}}$ $30.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $46.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.52}}$ $22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.38}}$
$\Delta$downarrow 6.5% $\downarrow$ 3.4% $\downarrow$ 30.7% $\downarrow$ 7.7% $\downarrow$ 6.0% $\downarrow$ 24.4%
@@
SD-20% 12 STONE $36.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ $53.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.44}}$ $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.08}}$ $41.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.23}}$ $63.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.11}}$ $37.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.29}}$
@@
w/ ST-TTC $35.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.08}}$ $52.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ $30.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ $40.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.41}}$ $60.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.07}}$ $33.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.73}}$
$\Delta$downarrow 2.9% $\downarrow$ 2.8% $\downarrow$ 10.6% $\downarrow$ 3.7% $\downarrow$ 3.8% $\downarrow$ 9.8%
@@
Avg. STONE $28.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.13}}$ $41.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.14}}$ $34.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.15}}$ $31.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.95}}$ $46.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.60}}$ $36.35{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.23}}$
@@
w/ ST-TTC $26.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.29}}$ $39.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ $24.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ $28.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.79}}$ $43.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.32}}$ $26.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±2.49}}$
$\Delta$downarrow 7.6% $\downarrow$ 4.8% $\downarrow$ 28.6% $\downarrow$ 8.0% $\downarrow$ 6.0% $\downarrow$ 26.8%
@@ E.2 Visualization Case
We provide more visualization examples of test set predictions to illustrate the effectiveness of our calibration, as shown in Figure 9
Figure 9: Show case of improvement in spatio-temporal forecasting through our ST-TTC .
@@ Appendix F More Discussion
F.1 Limitation
In this paper, we propose a novel paradigm for spatio-temporal forecasting: test-time computing. We systematically investigate the effective objectives for test-time computation and how to implement efficient computation. Based on detailed analysis, we posit that the key to the former lies in designing a lightweight, spatially-aware calibrator, while the key to the latter lies in designing an effective, non-information-leaking calibration parameter learning mechanism. Based on these insights, we present our corresponding solution, ST-TTC . We first propose a spectral domain calibrator, which appends a lightweight correction module operating in the frequency domain after the backbone network. This module calibrates errors by learning small, node-specific amplitude and phase correction factors. Furthermore, coupled with a flash gradient update mechanism featuring a streaming memory queue, it ensures universal, rapid, and resource-efficient test-time computing. Although there are still many potential areas for improvement, given the superiority and generality of our ST-TTC , we believe this provides a pathway for future exploration of larger-scale and more effective test-time computation. While we have taken a small step in this direction, several limitations warrant attention:

### v2

@@ E.1 Complete Results Table
w/ ST-TTC $34.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ $52.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.54}}$ $29.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.79}}$ $41.93{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.62}}$ $63.17{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$ $27.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.03}}$
$\Delta$ $\downarrow$ 1.3% $\downarrow$ 1.2% $\downarrow$ 12.5% $\downarrow$ 3.9% $\downarrow$ 4.4% $\downarrow$ 9.1%
@@
Avg. STONE $28.45{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $41.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.26}}$ $6.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.91}} 33$$3.20{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.69}}$ $49.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.97}}$ $29.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.00}}$
@@
w/ ST-TTC $26.59{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.22}}$ $39.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ $24.96{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.00}}$ $30.65{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$ $46.24{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.52}}$ $22.50{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.38}}$
$\Delta$ $\downarrow$ 6.5% $\downarrow$ 3.4% $\downarrow$ 30.7% $\downarrow$ 7.7% $\downarrow$ 6.0% $\downarrow$ 24.4%
@@
SD-20% 12 STONE $36.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.76}}$ $53.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.44}}$ $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.08}}$ $41.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.23}}$ $63.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.11}}$ $37.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.29}}$
@@
w/ ST-TTC $35.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.08}}$ $52.37{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.67}}$ $30.55{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.13}}$ $40.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.41}}$ $60.77{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.07}}$ $33.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.73}}$
$\Delta$ $\downarrow$ 2.9% $\downarrow$ 2.8% $\downarrow$ 10.6% $\downarrow$ 3.7% $\downarrow$ 3.8% $\downarrow$ 9.8%
@@
Avg. STONE $28.86{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.13}}$ $41.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.14}}$ $34.89{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.15}}$ $31.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.95}}$ $46.06{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.60}}$ $36.35{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 4.23}}$
@@
w/ ST-TTC $26.67{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.29}}$ $39.71{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $24.92{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$ $28.94{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.79}}$ $43.29{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.32}}$ $26.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 2.49}}$
$\Delta$ $\downarrow$ 7.6% $\downarrow$ 4.8% $\downarrow$ 28.6% $\downarrow$ 8.0% $\downarrow$ 6.0% $\downarrow$ 26.8%
@@ E.2 Visualization Case
We provide more visualization examples of test set predictions to illustrate the effectiveness of our calibration, as shown in Figure F.1
@@ Appendix F More Discussion
F.1 Distinction between Spatio-Temporal Forecasting and Long-Term Time Series
As stated in the paper, our work adheres to the common settings of previous short-term and long-term spatio-temporal forecasting studies (e.g., $12\rightarrow 12$, $24\rightarrow 24$) (Shao et al., 2022a, 2024b). However, we are fully aware of the $96\rightarrow 96-720$ settings prevalent in current long-term time series forecasting. We wish to share our insights regarding this:
❶There has been a long-standing debate concerning the long-term predictability of time series (Bergmeir, 2024; Brigato et al., 2025), which we do not intend to overly critique here. Nevertheless, it is important to note that in spatio-temporal forecasting, specifically in traffic flow theory research, previous studies (Zheng et al., 2025) utilizing residual analysis have indicated that, after minimizing periodicity, the correlation between traffic data beyond one hour and the past hour’s observations is significantly limited for most sensors. Furthermore, typical traffic sensor data collection frequency is approximately 5 minutes. Therefore, for practical traffic forecasting and decision-making scenarios, which usually focus on the next 1-2 hours (i.e., max 24 steps), our settings are more aligned with real-world applications.
❷Given that spatio-temporal forecasting can be considered a complex extension of time series forecasting, the additional dimension of sensor count (up to hundreds or thousands) results in an order of magnitude higher data training cost. This is a secondary reason why existing spatio-temporal forecasting often considers 12-step settings.
❸Another notable finding is that a recent study on the accuracy law of deep time series forecasting (Wang et al., 2025b), emerging subsequent to this paper, identifies a significant exponential relationship between the minimum prediction error of current deep forecasting models and the complexity of window series patterns. Adopting a Spectral-domain perspective similar to that of this paper, it defines the complexity of series patterns as the total variance of the amplitude spectrum distribution, thereby characterizing the intrinsic heterogeneity of series variations within each relevant window. This approach aligns with ours and provides a novel perspective to explain the learnability and effectiveness of the single-step gradient descent in our calibrator. While the study focuses solely on univariate time series forecasting, this limitation does not inherently undermine its relevance. It is worth noting that the study concludes that mainstream time series models have not yet reached saturation on traffic scenario-based benchmarks. However, we hold a different view, arguing that this conclusion primarily stems from the study’s exclusive use of time series forecasting models, while neglecting mainstream spatio-temporal dependency modeling models. For further in-depth discussions, we suggest that future research should address this aspect.
Figure F.1: Show case of improvement in spatio-temporal forecasting through our ST-TTC .
F.2 Limitation
In this paper, we propose a novel paradigm for spatio-temporal forecasting: test-time computing. Although there are still many potential areas for improvement, given the superiority and generality of our ST-TTC , we believe this provides a pathway for future exploration of larger-scale and more effective test-time computation. While we have taken a small step in this direction, several limitations warrant attention: